Evaluate the integral of te^(t²) dt


1
Expert's answer
2021-02-24T12:52:20-0500

"\int te^{t^2}dt=\frac{1}{2}\int e^{t^2} d(t^2)=\frac{1}{2} e^{t^2}+C"